OpenStudy (anonymous):

s^2-3s/s^2+s-12 (divided by) s+6/s+4? I need to factor it out. Here are the answer choices: A.s-3/s-6 B.s-3/s^2+6s C.s+6/s D. s/s+6

OpenStudy (dumbcow):

s^2 - 3s , both have an "s" in common, factor it out --> s(s-3) s^2 + s- 12, find 2 numbers that multiply to -12 but add to 1 --> (s + ?)(s + ?)

OpenStudy (mathstudent55):

Is this the problem? \(\dfrac{s^2 - 3s}{s^2 + s - 12} \div \dfrac{s + 6}{s + 4} \)

OpenStudy (anonymous):

Yes,That's the problem

OpenStudy (anonymous):

@dumbcow So,would the factors be -3,and 4?

OpenStudy (mathstudent55):

Factor the left numerator by simply factoring out s. Factor the left denominator like @dumcow showed above. Then replace the division sing with a multiplication sign and flip the second fraction.

OpenStudy (mathstudent55):

Correct, use -3 and 4.

OpenStudy (anonymous):

Could I factor the s^2-3s with 1 and 3?

OpenStudy (dumbcow):

s^2 - 3s , both have an "s" in common, factor it out --> s(s-3)

OpenStudy (mathstudent55):

No. That is simply factoring out an s as @dumbcow showed you earlier.

OpenStudy (mathstudent55):

After factoring the entire left fraction and flipping the right fraction and changing the division into a multiplication, you should have this: \(\dfrac{s(s - 3)}{(s + 4)(s - 3)} \times \dfrac{s + 4}{s + 6} \)

OpenStudy (anonymous):

Then the s+4's cancel out

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

then do the s-3's cancel too?

OpenStudy (anonymous):

then the answer is s/s+6?

OpenStudy (mathstudent55):

Yes. \(\dfrac{s\cancel{(s - 3)}}{\cancel{(s + 4)}\cancel{(s - 3)}} \times \dfrac{\cancel{s + 4}}{s + 6}\) \(= \dfrac{s}{s + 6} \)

OpenStudy (mathstudent55):

Good job!

OpenStudy (anonymous):

Okay,Thank you guys so much!

OpenStudy (mathstudent55):

You're welcome.